# PointMT: Efficient Point Cloud Analysis with Hybrid MLP-Transformer Architecture

### v1

@@ IV-G Neighborhood Size
IV-H Visual Analysis of Encoder Features
In contrast to traditional classification head, the Shape-Point Fusion (SPF) classification head fosters a more pronounced interaction between point features and encoded features during forward and backward propagation. As a result, the choice of classification heads can directly impact the encoded features. This section introduces two models: one with a traditional classification head and the other with an SPF classification head. To visually compare the feature representations of these models, t-SNE (t-Distributed Stochastic Neighbor Embedding) was employed to project the high-dimensional encoded features onto a two-dimensional plane. This facilitated a direct comparison of their distributional properties.
The t-SNE visualization exposes notable discrepancies in feature representation between the two models, as depicted in Fig. 4 and 5. When utilizing the SPF classification head, samples of the same class exhibited more tightly clustered patterns in the two-dimensional space, particularly the larger clusters with a higher sample count, which tended to adopt a more circular shape. This indicates improved precision in capturing within-class similarities and enhancing feature consistency. Additionally, the boundaries between clusters of different sizes corresponding to distinct classes were more clearly defined, with fewer overlapping regions. In contrast, the model with a traditional classification head demonstrated more ambiguous boundaries in certain areas of the encoded feature clustering distribution, indicating challenges in distinguishing between specific classes. The experimental results affirm the benefits of the SPF classification head in point cloud classification tasks. By effectively merging local point features with global pooled features, the SPF classification head not only enhances the model’s representational capacity for point cloud data but also enhances the clarity of class distinctions. This fusion approach equips the model with the ability to discern subtle features within point cloud data, thereby establishing a more structured feature space and achieving higher classification accuracy.
@@ IV-I1 Statistical Analysis of KL Divergence
The analysis indicates that when employing the SPF classification head, there is an increased alignment in the predictive outcomes of point features and pooled features, signifying improved feature consistency. This consistency enables point features to implicitly assimilate global information, thereby enhancing their discriminative capabilities. Simultaneously, the bolstered point features also augment the pooled features’ sensitivity to shape information, resulting in an overall enhancement in the classification model’s performance.

### v2

@@ IV-G Neighborhood Size
IV-H Visualization Analysis of Encoder Features and Logit Outputs
To verify the effectiveness of the SPF classification head, we conducted ablation experiments comparing it with a traditional classification head, using t-SNE (t-Distributed Stochastic Neighbor Embedding) for visual analysis of both encoded features and logit outputs. The results, shown in Fig. 4 and Fig. 5, are used to explore the differences in feature representation and classification performance between the two models.
As illustrated in Fig. 4, the t-SNE visualizations of the encoded features for the two models do not exhibit significant differences. This outcome can be attributed to the fact that these encoded features primarily represent the shape features obtained through pooling, without incorporating the point-level features that are fed into the SPF classification head and play a direct role in classification. Although quantitative experiments (Sec. LABEL:PointMT-sec-SPF) conducted later in this study demonstrated that the SPF classification head leads to notable differences in point features, these differences are not sufficiently pronounced to cause significant variation in the pooled shape features. Therefore, the t-SNE visualization of encoded features alone does not fully capture the impact of the classification heads on overall model performance.
The more meaningful comparison comes from the t-SNE visualization of the logit outputs, which directly reflect the classification outcomes, as shown in Fig. 5. In this figure, clear differences emerge between the two models. In Fig. 5(b), which corresponds to the SPF classification head, the clusters are more compact, and some exhibit a nearly circular distribution, indicative of higher classification confidence. Moreover, the clusters are more distinct, with clearer separations between them and fewer stray points, signifying more precise decision boundaries and improved generalization ability. These results demonstrate that the integration of shape and point features in the SPF classification head leads to enhanced classification performance, with greater confidence and superior generalization. Thus, the SPF classification head proves to be more effective in achieving robust and confident classification outcomes compared to the traditional approach.
@@ IV-I1 Statistical Analysis of KL Divergence
The analysis indicates that when employing the SPF classification head, there is an increased alignment in the predictive outcomes of point features and pooled features, signifying improved feature consistency. This consistency enables point features to implicitly assimilate global information, thereby enhancing their discriminative capabilities.